OpenStudy (anonymous):

solve for x: (x+1)^2 (x-2)+(x+1)(x-2)^2=0

myininaya (myininaya):

factor out (x-2)(x+1) (x-2)(x+1)[(x+1)+(x-2)]=0 (x-2)(x+1)(2x-1)=0

myininaya (myininaya):

x=2,x=-1,x=1/2